Given the following calculate the test statistic, If convenient, use technology:
a=0.10
Type of worker Professional Personal Both
Technical 38 35 34
Other 49 30 34

Answer by CPhill(2189)

To calculate the test statistic for this contingency table, we will use the chi-square test for independence.
1. Set up the Hypothesis:
Null Hypothesis (H0): The type of worker and the category (Professional, Personal, Both) are independent.
Alternative Hypothesis (H1): The type of worker and the category are dependent.
2. Create the Contingency Table:
Type of Worker Professional Personal Both Total
Technical 38 35 34 107
Other 49 30 34 113
Total 87 65 68 220
Export to Sheets
3. Calculate the Expected Frequencies:
The expected frequency for each cell is calculated as:
Expected Frequency = (Row Total * Column Total) / Grand Total
Expected (Technical, Professional) = (107 * 87) / 220 ≈ 42.34
Expected (Technical, Personal) = (107 * 65) / 220 ≈ 31.65
Expected (Technical, Both) = (107 * 68) / 220 ≈ 33.01
Expected (Other, Professional) = (113 * 87) / 220 ≈ 44.66
Expected (Other, Personal) = (113 * 65) / 220 ≈ 33.35
Expected (Other, Both) = (113 * 68) / 220 ≈ 34.99
4. Calculate the Chi-Square Statistic:
The chi-square statistic (χ²) is calculated using the formula:
χ² = Σ [(Observed - Expected)² / Expected]
χ² = [(38 - 42.34)² / 42.34] + [(35 - 31.65)² / 31.65] + [(34 - 33.01)² / 33.01] + [(49 - 44.66)² / 44.66] + [(30 - 33.35)² / 33.35] + [(34 - 34.99)² / 34.99]
χ² ≈ 0.457 + 0.354 + 0.029 + 0.413 + 0.334 + 0.028
χ² ≈ 1.615
Using Technology:
As shown by the provided python code, using the scipy library the chi-square statistic is calculated as 1.612983358368212.
Answer:
The test statistic (chi-square statistic) is approximately 1.613.